OpenStudy (anonymous):

Help!!?? What is the weekly payroll deduction if the deduction is 15%. Sharon gets paid $1,048 weekly? But I am told that the 15% is off the gross salary of $50,310. I only need to find the weekly payroll deductions.

OpenStudy (amistre64):

is the 1048 what she takes home? there seems to be some discrepency between the salary and the weekly pay

OpenStudy (amistre64):

1048*52=54496 so the weekly pay is greater than the yearly salary ... assuming that is what the salary defines

OpenStudy (littlebird):

Is the $50,310 the amount she gets payed in a year?

OpenStudy (debbieg):

If she takes home 1048 then 1048=0.85G where G=weekly gross pay. Then you can subtract G-1048=weekly deduction. But as @amistre64 points out, the other numbers aren't adding up.....

OpenStudy (debbieg):

If gross annual = 50,310, then annual net (after deduction) should be 0.85(50310) and weekly net should be 0.85(50310)/52=822.38, which differs from the 1048 you have above.

OpenStudy (anonymous):

Where is the .85 coming from? and yes that is accurate.

OpenStudy (debbieg):

If you deduct 15% from x, that is 0.15x. So what you have left is 0.85x. That is: x-(15% of x) = x-0.15x=x(1-0.15)=x(0.85) = 0.85x

OpenStudy (anonymous):

SO to find the answer would I need to divide $1,048 by .85?

OpenStudy (debbieg):

Well, that's what I originally tried, but again - the figures you are giving for net weekly salary and gross annual salary do not work out. So something is "off"... are you sure you're reading the problem correctly? Can you type out the exact problem as you have it?

OpenStudy (anonymous):

Yes, it is correct. Her yearly Salary is $50,310, her payroll deduction is 15% of her gross salary. She is estimated to get paid $1,048 weekly. And I need to know what her payroll deductions will be from her weekly pay.

OpenStudy (anonymous):

http://www.grossing.com/Salary-Calculator.htm this is the calculator that I used to find out her weekly pay.

OpenStudy (amistre64):

so you tried to calculate weekly pay? you can just deduct the yearly and divide by 52 for weekly

OpenStudy (debbieg):

OK, I thought that weekly pay amount was part of your given info in the problem. I'm not sure how that calculator is coming up with that weekly amount, because it is certainly NOT the annual/52.